# Low Complexity Channel Estimation for RIS-Assisted THz Systems with Beam Split

### v1

@@ III-A Angular Domain Sparse Representation
where $\mathbf{\Xi}_{\mathrm{R},m}\in\mathbb{C}^{N_{\mathrm{R}}\times Q_{\mathrm{B}}}$ is the simplified dictionary. Moreover, substituting $q_{2}=Q_{\mathrm{R}}$ into (14), the $q$-th column where $q=1,\dots Q_{\mathrm{B}}$ of $\mathbf{\Xi}_{\mathrm{R},m}$ can be formulated as
$\xi_{\mathrm{R},m,q}=\mathbf{a}_{\mathrm{R},m}\left(2\frac{q-Q_{\mathrm{R}}}{Q_{\mathrm{R}}}\right).$ (16)
By substituting the simplified $\mathbf{\Xi}_{\mathrm{R},m}$ into (13), the RIS-assisted cascaded channel can be rewritten as
@@
where $\mathbf{X}_{k,m}\in\mathbb{C}^{Q_{\mathrm{B}}\times Q_{\mathrm{T}}}$ is the total angle-excluded matrix, each row of $\mathbf{X}_{k,m}$ is a superposition of row subsets in $\bar{\mathbf{X}}_{k,m}\triangleq\bar{\boldsymbol{\beta}}_{k,m}\otimes\bar{\mathbf{\Upsilon}}_{m}$, i.e., $\mathbf{X}_{k,m}(i,:)=\sum_{j\in\mathcal{P}_{i}}{\bar{\mathbf{X}}_{k,m}}(j,:)$. $\mathcal{P}_{i}$ is the set of indices associated with the columns in $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ which are identical to the $i$-th column of $\mathbf{\Xi}_{\mathrm{R},m}$. Therefore, there are maximum $\Sigma_{k}=L\times J_{k}$ nonzero entries in $\mathbf{X}_{k,m}$. Employing the property of the Kronecker product, the vectorized RIS-assisted cascaded channel $\mathbf{h}_{k}^{\mathrm{cas}}[m]\in\mathbb{C}^{N_{\mathrm{T}}N_{\mathrm{R}}\times Q_{\mathrm{B}}Q_{\mathrm{T}}}$ is given by
@@
where $\mathbf{x}_{k,m}=\mathbf{X}_{k,m}(:)\in\mathbb{C}^{Q_{\mathrm{B}}Q_{\mathrm{T}}}$ represents $\mathbf{X}_{k,m}$ in vector form, $\mathbf{\Pi}_{m}\triangleq\left(\mathbf{C}_{\mathrm{T},m}^{*}\otimes\mathbf{\Xi}_{\mathrm{R},m}\right)\in\mathbb{C}^{N_{\mathrm{T}}N_{\mathrm{R}}\times Q_{\mathrm{B}}Q_{\mathrm{T}}}$ denotes the total dictionary. By substituting (18) into (3) and vectorizing, the sparse system formulation can be expressed as
@@
where $\mathbf{y}_{k,t}\left[m\right]\in\mathbb{C}^{P}$ is the vector form of $\tilde{\mathbf{y}}_{k,t}[m]$. Stacking the measurement vectors collected at $T$ successive subframes $\mathbf{y}_{k}\left[m\right]=\left[\mathbf{y}_{k,1}^{\mathsf{T}}\left[m\right],\cdots,\mathbf{y}_{k,T}^{\mathsf{T}}\left[m\right]\right]^{\mathsf{T}}\in\mathbb{C}^{PT}$, the received measurement signal can be expressed as
@@
where $\mathbf{\Psi}_{m}=\overline{\mathbf{R}}_{m}\mathbf{\Pi}_{m}$ and
$\overline{\mathbf{R}}_{m}\triangleq\left[\begin{array}[]{c}\mathbf{R}_{1}^{\mathsf{T}}[m]\otimes\boldsymbol{\theta}_{1}\\ \vdots\\ \mathbf{R}_{T}^{\mathsf{T}}[m]\otimes\boldsymbol{\theta}_{T}\\ \end{array}\right].$ (21)
It is obvious that the observation matrix $\mathbf{\Psi}_{m}\in\mathbb{C}^{PT\times Q_{\mathrm{B}}Q_{\mathrm{T}}}$ is a sparse matrix. So far, by employing the CBS dictionaries, the cascaded channel estimation problem has been transformed into the sparse signal recovery problem. Our primary task is to reconstruct the total angle-excluded vectors $\mathbf{x}_{k,m}$ with knowledge of the measurement vectors $\mathbf{y}_{k}[m]$ and the observation matrices $\mathbf{\Psi}_{m}$.
III-B Accurate Sparse Recovery and Channel Reconstruction
The complex sparse recovery problem has been formulated in Subsection III-A. To map this problem to the real number domain for resolution, we then transform the complex form in (20) to the equivalent real form
@@ III-B Accurate Sparse Recovery and Channel Reconstruction
By systematically assigning symbols to each matrix in (22), the sparse recovery problem is formulated as

### v2

@@ III-B2 Simplified Dictionary
where $\mathbf{\Xi}_{\mathrm{R},m}\in\mathbb{C}^{N_{\mathrm{R}}\times Q_{\mathrm{B}}}$ is the simplified dictionary.
Figure 3: Coupled dictionary simplification diagram.
Proof.
We first define the parameter sequence $\bar{\vartheta}_{q_{1}}=\delta\cdot q_{1}-(Q+1)/{Q},\forall q_{1}=1,\dots,Q_{\mathrm{R}}$ which corresponds one-to-one with $Q_{\mathrm{R}}$ ARV columns in the dictionary $\mathbf{C}_{\mathrm{R},m}$. Similarly, we define the sequence $\delta_{q_{2}}=-\delta\cdot q_{2}+(Q+1)/Q$ to correspond to $Q_{R}$ ARV columns in the conjugate dictionary $\mathbf{C}_{\mathrm{R},m}^{*}$. Thus, the column vectors of $\mathbf{C}_{\mathrm{R},m}^{*}\bullet\mathbf{C}_{\mathrm{R},m}$ can be expressed as $\mathbf{a}_{\mathrm{R},m}(\bar{\vartheta}_{q_{1}}-\bar{\vartheta}_{q_{2}})$ for $q_{1},q_{2}=1,\dots,Q_{\mathrm{R}}$. The $j$-th column of the coupled dictionary $\tilde{\boldsymbol{\xi}}_{\mathrm{R},m,j}$ can be expressed as
$\tilde{\boldsymbol{\xi}}_{\mathrm{R},m,j}=\mathbf{a}_{\mathrm{R},m}(\delta\left(q_{1}-q_{2}\right)),$ (16)
where $j=\left(q_{2}-1\right)Q_{\mathrm{R}}+q_{1},\forall q_{1},q_{2}=1,\dots,Q_{\mathrm{R}}$. Based on the fact that $q_{1}-q_{2}\in\left[1-Q_{\mathrm{R}},Q_{\mathrm{R}}-1\right]$, the number of unique columns of coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ is $Q_{\mathrm{B}}=2Q_{\mathrm{R}}-1$.
Then, we simplify the coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ step by step from (a) to (d) as shown in Fig. 3. As depicted in step (a), we reshape the overcomplete dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ of length $Q_{\mathrm{R}}^{2}$ ARV columns into a three-dimensional matrix where each column represents an ARV. Given that ARV vectors can be uniquely determined by known angles in (16), we use the sequences $q_{1},q_{2}=1,\dots,Q_{\mathrm{R}}$ to index the corresponding ARV columns. In step (b), we form the index matrix $\mathbf{Q}\in\mathbb{N}^{Q_{\mathrm{R}}\times Q_{\mathrm{R}}}$ with elements $[\mathbf{Q}]_{i,j}=q_{1}(i)-q_{2}(j)$ for $i,j=1,\dots,Q_{\mathrm{R}}$. It is noticed that due to the cyclic shift property of this dictionary, the dictionary base can be represented entirely by combining just the first and last rows. This means that we can reconstruct the original dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ with a sliding window of length $Q_{\mathrm{R}}$. Therefore, the simplified CBS dictionary $\mathbf{\Xi}_{\mathrm{R},m}$ can be given as in (15). Moreover, substituting $q_{2}=Q_{\mathrm{R}}$ into (16), the $q$-th column where $q=1,\dots Q_{\mathrm{B}}$ of $\mathbf{\Xi}_{\mathrm{R},m}$ can be formulated as
$\boldsymbol{\xi}_{\mathrm{R},m,q}=\mathbf{a}_{\mathrm{R},m}\left(2\frac{q-Q_{\mathrm{R}}}{Q_{\mathrm{R}}}\right).$ (17)
III-B3 Sparse Recovery Reformulation
By substituting the simplified CBS dictionary $\mathbf{\Xi}_{\mathrm{R},m}$ into (13), the RIS-assisted cascaded channel can be rewritten as
@@ III-B3 Sparse Recovery Reformulation
where $\mathbf{X}_{k,m}\in\mathbb{C}^{Q_{\mathrm{B}}\times Q_{\mathrm{T}}}$ is the total angle-excluded matrix, each row of $\mathbf{X}_{k,m}$ is a superposition of row subsets in $\bar{\mathbf{X}}_{k,m}\triangleq\bar{\boldsymbol{\beta}}_{k,m}\otimes\bar{\mathbf{\Sigma}}_{m}$, i.e., $\mathbf{X}_{k,m}(i,:)=\sum_{j\in\mathcal{P}_{i}}{\bar{\mathbf{X}}_{k,m}}(j,:)$. $\mathcal{P}_{i}$ denotes the set of indices associated with the columns in $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ that match the $i$-th column of $\mathbf{\Xi}_{\mathrm{R},m}$. Therefore, there are at most ${L}_{\mathrm{tot},k}=L\times J_{k}$ nonzero entries in $\mathbf{X}_{k,m}$. Employing the Kronecker product property, the vectorized RIS-assisted cascaded channel $\mathbf{h}_{k}^{\mathrm{cas}}[m]\in\mathbb{C}^{N_{\mathrm{T}}N_{\mathrm{R}}\times Q_{\mathrm{B}}Q_{\mathrm{T}}}$ can be given by
@@
where $\mathbf{x}_{k,m}=\mathbf{X}_{k,m}(:)\in\mathbb{C}^{Q_{\mathrm{B}}Q_{\mathrm{T}}}$ represents $\mathbf{X}_{k,m}$ in vector form, $\mathbf{\Pi}_{m}\triangleq\left(\mathbf{C}_{\mathrm{T},m}^{*}\otimes\mathbf{\Xi}_{\mathrm{R},m}\right)\in\mathbb{C}^{N_{\mathrm{T}}N_{\mathrm{R}}\times Q_{\mathrm{B}}Q_{\mathrm{T}}}$ denotes the total dictionary.
By substituting (19) into (2) and vectorizing, the sparse system formulation can be expressed as
@@
where $\mathbf{y}_{k,t}\left[m\right]\in\mathbb{C}^{P}$ is the vector form of $\tilde{\mathbf{y}}_{k,t}[m]$. Stacking the measurement vectors collected at $T$ successive subframes $\mathbf{y}_{k}\left[m\right]=[\mathbf{y}_{k,1}^{\mathsf{T}}\left[m\right],\cdots,\mathbf{y}_{k,T}^{\mathsf{T}}\left[m\right]]^{\mathsf{T}}\in\mathbb{C}^{PT}$, the received signal can be expressed as
@@
where $\mathbf{\Psi}_{m}=\bar{\mathbf{R}}_{m}\mathbf{\Pi}_{m}$ and
$\bar{\mathbf{R}}_{m}\triangleq\left[\begin{array}[]{c}\mathbf{R}_{1}^{\mathsf{T}}[m]\otimes\boldsymbol{\theta}_{1}\\ \vdots\\ \mathbf{R}_{T}^{\mathsf{T}}[m]\otimes\boldsymbol{\theta}_{T}\\ \end{array}\right].$ (22)
Note that the observation matrix $\mathbf{\Psi}_{m}\in\mathbb{C}^{PT\times Q_{\mathrm{B}}Q_{\mathrm{T}}}$ is a sparse matrix. To map this problem to the real number domain for resolution, we then transform the complex form in (21) to the equivalent real form
@@
By systematically assigning symbols to each matrix in (23), the sparse recovery problem can be reformulated as